# Spin-triplet pair density wave superconductors

### v1

@@ Abstract
Recent experiments have shown that the nonzero center of mass momentum pair density wave (PDW) is a widespread phenomenon observed over different superconducting materials. However, concrete theoretical model realizations of the PDW order have remained elusive. Here, we study a one-dimensional model with nearest-neighbor pairing attraction, i.e. a spinful Kitaev chain, under generic spin-orbit couplings such that the spin-rotation symmetry is fully broken. The most general superconducting order parameter is described by a spatial dependent $\mathbf{d}_{i}$-vector. We show that a spin-triplet pair density wave (t-PDW) emerges in the ground state and occupies a large part of the phase diagram. The $\mathbf{d_{i}}$-vector of the t-PDW rotates with a pitch $Q_{\rm pdw}$ along the chain and spans an ellipsoid. The pure t-PDW is fully-gapped and a class-DIII topological superconductor with two Majorana zero modes localized at each end of the chain and protected by time-reversal symmetry. Our findings reveal unprecedented insights into the exotic pure PDW superconductor and provide a possible explanation for the one-dimensional PDW detected along domain walls in monolayer iron-based superconductor Fe(Te,Se) and potentially realizable using other quantum structures in unconventional superconductors.
@@
The FFLO states can be further divided into FF and LL states. The Cooper pairs in the FF state carry a single nonzero momentum $\mathbf{q}$, such that the SC order parameter breaks time-reversal symmetry and varies in space as $\Delta(\mathbf{r})=\Delta_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}}$. This helical superconductor has been extensively studied in noncentrosymmetric systems with spin-orbit coupling (SOC) and time-reversal symmetry-breaking fields [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. It is believed to play an important role in realizing the diode effect in SC systems [18, 19, 20]. The LO state, on the other hand, is described by a SC order parameter containing at least two nonzero momenta i.e., ±$\pm\mathbf{q}$, and varies in space according to $\Delta(\mathbf{r})=\Delta_{\mathbf{q}}\cos{(\mathbf{q}\cdot\mathbf{r})}$. This inhomogeneous SC state does not require broken time-reversal symmetry. Its generalization in the absence of an external magnetic field is more commonly referred to as the pair density wave (PDW) state [21].
The PDW state has been proposed to exist in the high-$T_{c}$ cuprate superconductors [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and has attracted intense research interest [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]. More recently, evidence for PDW formation has been reported in iron-based superconductors [71, 72], heavy-fermion superconductors [73, 74], and kagome superconductors [75, 76]. Despite these exciting developments, concrete theoretical model realizations of the PDW states beyond phenomenological Landau free-energy descriptions have remained very challenging.
Motivated by the recent experimental progress, especially the detection of the one-dimensional PDW modulations at the domain walls in the iron-based superconductor Fe(Te,Se) film [72], we propose and study a one-dimensional (1D) model with nearest-neighbor pairing attraction and spin-orbit couplings (SOCs) such that the spin-rotation symmetry is fully broken. This simple and concrete model can be thought as a spinful Kitaev chain in the presence of both Rashba ($\alpha_{R}$) and Dresselhaus ($\alpha_{D}$) SOC. Since the spin rotation symmetry is fully broken, the pairing interaction can be parametrized independently by attractions in the equal-spin ($V_{1}$) and opposite-spin ($V_{2}$) channels. The nature of the SC states can thus be explored by varying the ratio of $r_{V}=V_{2}/V_{1}$ and $r_{\rm so}=\alpha_{D}/\alpha_{R}$.
Due to the broken inversion symmetry, the pairing state is in general of a mixed parity type, with the spin-triplet sector specified by the $d$-vector [77]. The most general SC state can thus be described by a spatial-dependent pairing order parameter (Methods)
@@
where $\psi_{i}$ and $\mathbf{d}_{i}$ correspond to the spin-singlet and the spin-triplet components at site-$i$, respectively. We perform self-consistent mean-field theory calculations and explore the phase diagram as a function of the SOC ratio $r_{\rm so}$ and the ratio of opposite and equal-spin attractions $r_{V}$. We find that a time-reversal symmetric, spin-triplet pair density wave (t-PDW) superconductor emerges in the ground state and occupies a large part of the phase diagram, which also contains a uniform mixed-parity (MP) superconductor. For a fixed SOC ratio $r_{\rm so}$, the phase diagram is shown schematically shown in Fig. 1, displaying a transition from the uniform MP topological superconductor (TSC) for $V_{2}>V_{1}$ to a t-PDW TSC for $V_{1}>V_{2}$. In the t-PDW phase, the spin-singlet component vanishes, while the $\mathbf{d_{i}}$-vector rotates with a pitch $Q_{\rm pdw}$ along the chain and spans an ellipsoid (Fig. 1). We will show that this pure t-PDW is fully-gapped and belongs to the class-DIII TSC with two Majorana zero modes localized at each end of the chain and protected by time-reversal symmetry. Moreover, we study the evolution of the phase diagram with the chemical potential and the properties of the topological phase transitions. Our simple and concrete model allows us to address the critical question of how the pairing interaction and SOC work together to overcome the kinetic energy of the pairs with nonzero center of mass momentum, such that the t-PDW can emerge in the ground state.
@@ Results
This effective Hamiltonian can describe the embedded quantum structures in high-Tc superconductors due to spatial symmetry breaking [78], such as along the atomic line defects in monolayer Fe(Te,Se) [79]. The effective attraction $V_{1}$ and $V_{2}$ can be different since the system already breaks the spin rotation symmetry. Here, the interaction term breaks SU(2) spin rotation symmetry down to U(1) symmetry about $z$ axis when $V_{1}\neq V_{2}$, while the SOC term breaks the SU(2) symmetry down to U(1) symmetry about a general axis determined by $\alpha_{R}$ and $\alpha_{D}$, therefore the Hamiltonian fully breaks the spin rotation symmetry, which is important to realize the PDW state. The two attraction terms can be decomposed into equal-spin and opposite-spin pairing channels so that the full Hamiltonian can be written as
@@
and the non-interacting Hamiltonian is written as
@@
Figure 1: Triplet PDW state. Schematic phase diagram as the ratio of the attraction $r_{V}=\frac{V_{2}}{V_{1}}$ is tuned for a small value of the ratio of spin-orbit coupling $r_{\rm so}=\frac{\alpha_{D}}{\alpha_{R}}$. The phase transition between the t-PDW state and the uniform mixed-parity TSC (MPTSC) state occurs at the transition point with $V_{1}=V_{2}$. The t-PDW state is a pure spin-triplet pairing state that can be described by a spatial-dependent imaginary $d$-vector. Its imaginary part is depicted by red arrows showing the spatial modulation. The $d$-vector revolves around the origin in an elliptical orbit in 3D space within a period of the t-PDW state. The uniform order parameters describe the MPTSC, i.e. both $\psi$ and d are uniform in real space.
@@
Here, the electrons belonging to the opposite branches can form Cooper pairs with zero center of mass momentum. Since each branch of the bands contains both spin components, the pairing can be either equal-spin pairing $\Delta_{\parallel,0}$ or opposite-spin pairing $\Delta_{\perp,0}$. While the equal-spin pairing must be spin-triplet, the opposite-spin pairing is in general mixed-parity pairing where the mixture of the spin-singlet and triplet pairing is controlled by the phase of the order parameter $\Delta_{\perp,0}$, with phase 00 corresponding to the $s$-wave pairing and phase $\pi$/2 corresponding to the $p$-wave pairing and other phase values corresponding to the mixture of both. On the other hand, the electrons belonging to the same branch can pair up with the nonzero center of mass momentum $Q$ and $-Q$, which is determined by the SOC as $Q=2\arctan\left(\frac{\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}}{t}\right)$. Due to the same reason, the finite momentum intra-branch pairing can also be in the equal-spin ($\Delta_{\parallel,Q}$, $\Delta_{\parallel,-Q}$) and opposite-spin pairing channels ($\Delta_{\perp,Q}$, $\Delta_{\perp,-Q}$). The mean-field equations are then solved with a fixed chemical potential $\mu$ in the folded Brillouin Zone (BZ) for the commensurate $Q$. Assuming $Q=\frac{\pi}{D}$, the BZ is folded into $\frac{1}{2D}$ of the original BZ.
Figure 2: Schematic band structure of the spin-orbit coupled chain, showing different pairing channels across the Fermi points. a The inter-branch pairing leads to the uniform pairing order with zero center of mass momentum for the Cooper pair q=0 ($\Delta_{0}$). b The intra-branch pairing gives rise to the pairing states with the finite center of mass momentum q=±$\pm Q$ ($\Delta_{\pm Q}$). Both pairings can be equal-spin and opposite-spin states.
Full Phase diagram
@@
We first choose the parameters as $t=1$, $\mu=-0.4$, $\sqrt{V_{1}^{2}+V_{2}^{2}}=1.5$, $\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}=\tan\left(\pi/8\right)$, so that $Q=\frac{\pi}{4}$ and tune the ratio $r_{\rm so}$ and $r_{V}$. The ground state should be determined by comparing the energy of possible mean-field solutions and we find that the two most competitive states are the t-PDW state and the uniform MP state. The obtained phase diagram as functions of $r_{\rm so}$ and $r_{V}$ are shown in Fig. 3, where two phase boundaries separate the t-PDW and MP states and cross at a single point. The phase diagram has several features. First of all, when $r_{\rm so}=0$, the ground state is always the Kramers Fulde-Ferrell (KFF) state for $r_{V}<1$ until it reaches the transition point at $r_{V}=1$ and the system transits to the MP state through a first-order phase transition which is studied in the previous work [20]. The KFF can be viewed as a special limit of the t-PDW state, where the Cooper pairs carry opposite momenta in the opposite spin channels. It has no spatial modulation in the local density of states, although its pairing function carries the finite center of mass momenta. This can be proved by connecting the KFF state to a uniform $p$-wave superconductor through a gauge transformation [20]. For the case with $r_{V}<1$, introducing finite $r_{\rm so}$ invalidates such gauge transformation so that the ground state becomes the t-PDW state where the spatial modulation of the order parameters can no longer be gauged away. As $r_{\rm so}$ further increases, the system goes through a first-order phase transition and the ground state becomes the uniform MP state. Moreover, for the case with $r_{V}>1$, the ground state remains the uniform MP state when the finite $r_{\rm so}$ is introduced until the ratio is big enough to induce a first-order transition from the MP state to the t-PDW state. While the horizontal transition line is determined numerically, the vertical transition line can be determined analytically and is verified in the numerical calculation.
@@
Figure 3: Phase diagrams close to half filling. The t-PDW and MPTSC states are realized by tuning the two ratios $r_{\rm so}$ and $r_{V}$ in the mean-field calculations, where the chemical potential $\mu=-0.4$ close to half filling. The KFF state is realized in the $x$-axis with $r_{\rm so}=0$ and $r_{V}<1$ represented by the bold blue line.
@@
This vertical transition line corresponds to the case with $r_{V}=1$, where the interaction part of the Hamiltonian has the full SU(2) spin rotation symmetry. Then for the 1D system, the SOC reduces this symmetry to the U(1) symmetry corresponding to the spin rotation around the special axis with unit vector $\hat{n}_{\alpha}=(\alpha_{D},0,\alpha_{R})/\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}$. Then we can always choose the spin quantization axis along the axis $\hat{n}_{\alpha}$ and the Hamiltonian thus becomes
@@
Figure 4: Properties of typical t-PDW and MP states shown in the phase diagram of Fig. 3. a,c The band dispersion of t-PDW (a) and MP (c) states, where that of the t-PDW state is plotted in the folded Brillouin zone. b,d The energy spectrum obtained in a finite-size system with L=320, hosting four zero-energy MZMs denoted by blue and red dots. The parameters used are $r_{\rm so}$=1.4 and $r_{V}$=0.3 for the t-PDW state and $r_{\rm so}$=3.0 and $r_{V}$=0.3 for the MP state with $\mu=-0.4$ and all other parameters mentioned in the main text.
Real space $d$-vectors for t-PDW states
We next study the solved pairing order parameters of all the states in the phase diagram, which helps to visualize the spatial modulation of the t-PDW state. As mentioned above, the real space pairing order parameters can be compactly expressed by Eq. 1 as
@@
The real space $d$-vector is pure imaginary due to the time-reversal symmetry. The spatial modulation of a typical $d$-vector of the t-PDW state constructed from the pairing order parameters determined self-consistently with parameters $r_{\rm so}=1.4$ and $r_{V}=0.3$ is shown schematically in Fig. 1, where the $d$-vector revolves around the origin in an elliptical orbit in 3D space within a period of the PDW state.
Moreover, the KFF state got for $r_{\rm so}=0$ and $r_{V}<1$ has only one finite order parameter $\Delta_{\parallel,Q}$, corresponding to the $d$-vector as
@@
The band dispersion of both t-PDW and MP states are shown in Fig. 4(a,c), both of which are fully gapped. In order to determine the topological properties of these states, we also calculate the end states of the finite-size systems for each phase. As shown in Fig. 4(b,d), both phases can host two pairs of Majorana zero modes (MZMs) indicating the nontrivial topological properties for both phases. Since the time-reversal symmetry is preserved for both phases, they can be classified as time-reversal invariant topological superconductors. The nontrivial topological property of the t-PDW state is easy to understand since it is a pure $p$-wave spin-triplet pairing state with $d$-vector given by Eq. 8 and the spatial modulation of the $d$-vector does not change its topological property. While for the MP state with both $s$-wave and $p$-wave pairing components, the state can be either topological superconductor or trivial superconductor [78]. For the chemical potential close to half-filling with $\mu=-0.4$, the MP state is always topological, which is labeled as MPTSC in the phase diagram of Fig. 3. However, a topological phase transition can occur when the chemical potential is tuned away from the half-filling, which is studied and discussed below.
@@
Finally, we also study the effect of the chemical potential on the phase diagram. The calculation above is done for $\mu=-0.4$ which is around the half-filling of the chain. We find that if the chemical potential is tuned away from the half-filling, the region for the t-PDW phase shrinks as shown in Fig. 5a with $\mu=-1.4$, and for $\mu$ close enough to the band top or bottom, the phase diagram changes dramatically, where the t-PDW state can only be realized for small values of $r_{V}$ and $r_{\rm so}$ which is shown in Fig. 5b with $\mu=-1.8$.
Figure 5: Phase diagrams away from half filling. The chemical potential is set as $\mu$=-1.4 (a) and $\mu=-1.8$ (b) with all the other parameters unchanged. The KFF state is realized in the $x$-axis labeled by the bold blue line. The red lines represent the transition line between the TSC and trivial SC within the MP phase.
@@
where $t_{\alpha}=\sqrt{t^{2}+\alpha_{R}^{2}+\alpha_{D}^{2}}$, $\theta_{\alpha}=\arctan\left(\frac{\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}}{t}\right)=\frac{Q}{2}$ and $n^{d}_{i\sigma}=d_{i\sigma}^{\dagger}d_{i\sigma}$. Since the SOC is gauged away in the transformed Hamiltonian, which respects the full spin rotation and inversion symmetry, the ground state of such a system is either the uniform $s$-wave pairing state or the $p$-wave state with its $d$-vector pointing in any direction, depending on the chemical potential. For $\mu$ around the half filling, the Fermi points $k_{f}$ are round ±$\pm\frac{\pi}{2}$ where the pairing function for the $p$-wave state $\Delta_{p}(k)\propto\sin k$ gains more energy than that of the $s$-wave state $\Delta_{s}(k)\propto\cos k$, making the $p$-wave pairing state as the ground state. On the other hand, if $\mu$ is sufficiently tuned away from the half-filling, the $s$-wave pairing will take over as $k_{f}$ approaches ±$\pm\pi$, where $\Delta_{s}(k)\propto\cos k$ gains more energy.

### v2

@@ Abstract
Recent experiments have shown that the nonzero center of mass momentum pair density wave (PDW) is a widespread phenomenon observed over different superconducting materials. However, concrete theoretical model realizations of the PDW order have remained elusive. Here, we study a one-dimensional model with nearest-neighbor pairing attraction, i.e. a spinful Kitaev chain, under generic spin-orbit couplings such that the spin-rotation symmetry is fully broken. The most general superconducting order parameter is described by a spatial dependent $\mathbf{d}_{i}$-vector. We show that a spin-triplet pair density wave (t-PDW) emerges in the ground state and occupies a large part of the phase diagram. The $\mathbf{d_{i}}$-vector of the t-PDW rotates with a pitch $Q_{\rm pdw}$ along the chain and spans an ellipsoid. The pure t-PDW is fully gapped and realizes a class-DIII topological superconductor, characterized by two Majorana zero modes localized at each end of the chain and protected by time-reversal symmetry. Our findings reveal unprecedented insights into the exotic pure PDW superconductor and provide a possible explanation for the one-dimensional PDW detected along domain walls in monolayer iron-based superconductor Fe(Te,Se) and potentially realizable using other quantum structures in unconventional superconductors.
Introduction
@@
The FFLO states can be further divided into FF and LL states. The Cooper pairs in the FF state carry a single nonzero momentum $\mathbf{q}$, such that the SC order parameter breaks time-reversal symmetry and varies in space as $\Delta(\mathbf{r})=\Delta_{\mathbf{q}}e^{i\mathbf{q}\cdot\mathbf{r}}$. This helical superconductor has been extensively studied in noncentrosymmetric systems with spin-orbit coupling (SOC) and time-reversal symmetry-breaking (TRSB) fields [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. It is believed to play an important role in realizing the diode effect in SC systems [18, 19, 20]. The LO state, on the other hand, is described by a SC order parameter containing at least two nonzero momenta i.e., $\pm\mathbf{q}$, and varies in space according to $\Delta(\mathbf{r})=\Delta_{\mathbf{q}}\cos{(\mathbf{q}\cdot\mathbf{r})}$. This inhomogeneous SC state does not require broken time-reversal symmetry. Its generalization in the absence of an external magnetic field is more commonly referred to as the pair density wave (PDW) state [21].
The PDW state has been proposed to exist in the high-$T_{c}$ cuprate superconductors [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and the has attracted intense research interest [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. Moreover, the PDW state is also studied in the spin-triplet pairing phase of He-3 superfluid [75, 76, 77, 78, 79, 80]. More recently, evidence for PDW formation has been reported in iron-based superconductors [81, 82], heavy-fermion superconductors [83, 84], and kagome superconductors [85, 86]. Despite these exciting developments, concrete theoretical model realizations of the PDW states beyond phenomenological Landau free-energy descriptions have remained very challenging.
Motivated by the recent experimental progress, especially the detection of the one-dimensional PDW modulations at the domain walls in the iron-based superconductor Fe(Te,Se) film [82], we propose and study a one-dimensional (1D) model with nearest-neighbor pairing attraction and SOCs such that the spin-rotation symmetry is fully broken. This simple and concrete model can be thought as a generalization of the spinless Kitaev model to the spinful case in the presence of both Rashba ($\alpha_{R}$) and Dresselhaus ($\alpha_{D}$) SOC which couples the two spin species and maintains time-reversal symmetry. Since the spin rotation symmetry is fully broken, the pairing interaction can be parametrized independently by attractions in the equal-spin ($V_{1}$) and opposite-spin ($V_{2}$) channels. The nature of the SC states can thus be explored by varying the ratio of $r_{V}=V_{2}/V_{1}$ and $r_{\rm so}=\alpha_{D}/\alpha_{R}$.
Due to the broken inversion symmetry, the pairing state is in general of a mixed parity type [87], with the spin-triplet sector specified by the $d$-vector [88]. The most general SC state can thus be described by a spatial-dependent pairing order parameter (Methods)
@@
where $\psi_{i}$ and $\mathbf{d}_{i}$ correspond to the spin-singlet and the spin-triplet components at site-$i$, respectively. We perform self-consistent mean-field theory calculations and explore the phase diagram as a function of the SOC ratio $r_{\rm so}$ and the ratio of opposite and equal-spin attractions $r_{V}$. We find that a time-reversal symmetric, spin-triplet pair density wave (t-PDW) superconductor emerges in the ground state and occupies a large part of the phase diagram, which also contains a uniform mixed-parity (MP) superconductor. For a fixed SOC ratio $r_{\rm so}$, the phase diagram is shown schematically shown in Fig. 1, displaying a transition from the uniform MP topological superconductor (TSC) for $V_{2}>V_{1}$ to a t-PDW TSC for $V_{1}>V_{2}$. In the t-PDW phase, the spin-singlet component vanishes, while the $\mathbf{d_{i}}$-vector rotates with a pitch $Q_{\rm pdw}$ along the chain and spans an ellipsoid (Fig. 1). We will show that this pure t-PDW is fully-gapped and belongs to the class-DIII TSC with two Majorana zero modes localized at each end of the chain and protected by time-reversal symmetry. Moreover, we study the evolution of the phase diagram with the chemical potential and the properties of the topological phase transitions. Our simple and concrete model allows us to address the critical question of how the pairing interaction and SOC work together to overcome the kinetic energy of the pairs with nonzero center of mass momentum, such that the t-PDW can emerge in the ground state. The mean-field theory is formulated at zero temperature, where quasi-long-range order can exist with power-law correlations in 1D quantum systems when fluctuations are included beyond the mean-field theory.
@@ Results
This effective Hamiltonian can describe the embedded quantum structures [89], such as the domain wall [82] and atomic line defects [90] in monolayer Fe(Te,Se) grown on SrTiO3. The effective attraction $V_{1}$ and $V_{2}$ can be different since the system already breaks the spin rotation symmetry. Here, the interaction term breaks SU(2) spin rotation symmetry down to U(1) about $z$ axis when $V_{1}\neq V_{2}$, while the SOC term breaks the SU(2) symmetry down to U(1) about a general axis determined by $\alpha_{R}$ and $\alpha_{D}$, therefore the Hamiltonian fully breaks the spin rotation symmetry. The connection between the model Hamiltonian and the 1D quantum structure in monolayer Fe(Te,Se) stems from their symmetry properties. In particular, the inversion symmetry breaking due to the substrate induces the Rashba SOC and the mirror symmetry breaking along the domain wall or line defects introduces the Dresselhaus SOC. As a result, the spin-rotation symmetry is completely broken, which we show plays a crucial role in the realization of the PDW state. The two attraction terms can be decomposed into equal-spin and opposite-spin pairing channels so that the full Hamiltonian can be written as
@@
and the non-interacting Hamiltonian has the form
@@
Figure 1: Spin-triplet pair density wave (t-PDW) state. Schematic phase diagram as the ratio of the attraction $r_{V}=\frac{V_{2}}{V_{1}}$ is tuned for a small value of the ratio of spin-orbit coupling $r_{\rm so}=\frac{\alpha_{D}}{\alpha_{R}}$. The phase transition between the t-PDW state and the uniform mixed-parity topological superconductor (MPTSC) state occurs at the transition point with $V_{1}=V_{2}$. The t-PDW state is a pure spin-triplet pairing state that can be described by a spatial-dependent imaginary $d$-vector. Its imaginary part is depicted by red arrows showing the spatial modulation. The $d$-vector revolves around the origin in an elliptical orbit in 3D space within a period of the t-PDW state. The uniform order parameters describe the MPTSC, i.e. both $\psi$ and d are uniform in real space.
@@
Here, the electrons belonging to the opposite branches can form Cooper pairs with zero center of mass momentum. Since each branch of the bands contains both spin components, the pairing can be either equal-spin pairing $\Delta_{\parallel,0}$ or opposite-spin pairing $\Delta_{\perp,0}$. While the equal-spin pairing must be spin-triplet, the opposite-spin pairing is in general mixed-parity pairing where the mixture of the spin-singlet and triplet pairing is controlled by the phase of the order parameter $\Delta_{\perp,0}$, with phase 0 corresponding to the $s$-wave pairing and phase $\pi$/2 corresponding to the $p$-wave pairing and other phase values corresponding to the mixture of both. On the other hand, the electrons belonging to the same branch can pair up with the nonzero center of mass momentum $Q$ and Q-Q, which is determined by the SOC as $Q=2\arctan\left(\frac{\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}}{t}\right)$. Due to the same reason, the finite momentum intra-branch pairing can also be in the equal-spin ($\Delta_{\parallel,Q}$, $\Delta_{\parallel,-Q}$) and opposite-spin pairing channels ($\Delta_{\perp,Q}$, $\Delta_{\perp,-Q}$). The mean-field equations are then solved with a fixed chemical potential $\mu$ in the folded Brillouin Zone (BZ) for the commensurate $Q$. Assuming $Q=\frac{\pi}{D}$, the BZ is folded into $\frac{1}{2D}$ of the original BZ.
[…]
Phase diagram
@@
We first choose the parameters as $t=1$, $\mu=-0.4$, $\sqrt{V_{1}^{2}+V_{2}^{2}}=1.5$, $\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}=\tan\left(\pi/8\right)$, so that $Q=\frac{\pi}{4}$ and tune the ratio $r_{\rm so}$ and $r_{V}$. The ground state should be determined by comparing the energy of possible mean-field solutions and we find that the two most competitive states are the t-PDW state and the uniform MP state. The obtained phase diagram as functions of $r_{\rm so}$ and $r_{V}$ is shown in Fig. 3, where two phase boundaries separate the t-PDW and MP states and cross at a single point. The phase diagram has several features. First of all, when $r_{\rm so}=0$, the ground state is always the Kramers Fulde-Ferrell (KFF) state for $r_{V}<1$ until it reaches the transition point at $r_{V}=1$ and the system transits to the MP state through a first-order phase transition which is studied in the previous work [20]. The KFF can be viewed as a special limit of the t-PDW state, where the Cooper pairs carry opposite momenta in the opposite spin channels. It has no spatial modulation in the local density of states, although its pairing function carries the finite center of mass momenta. This can be proved by connecting the KFF state to a uniform $p$-wave superconductor through a gauge transformation [20]. However, when $r_{\rm so}$ is nonzero, the ground state becomes the pure t-PDW state for $r_{\rm so}<1$ since the spatial modulation of the order parameter can no longer be gauged away. As $r_{\rm so}$ further increases, the system goes through a first-order phase transition and the ground state becomes the uniform MP state. Intriguingly, the phase structure is inverted for the case where $r_{V}>1$. The ground state remains the uniform MP state when nonzero $r_{\rm so}$ is introduced until the ratio is big enough to induce a first-order transition from the MP state to the t-PDW state. While the horizontal transition line is determined numerically, the vertical transition line can be determined analytically and is verified by numerical calculations.
@@
Figure 3: Phase diagrams close to half filling. The spin-triplet pair density wave (t-PDW) and the mixed-parity topological superconductor (MPTSC) states are realized by tuning the two ratios $r_{\rm so}$ and $r_{V}$ in the mean-field calculations, where the chemical potential $\mu=-0.4$ close to half filling. The Kramers Fulde-Ferrell (KFF) state is realized in the $x$-axis with $r_{\rm so}=0$ and $r_{V}<1$ represented by the bold blue line.
@@
The vertical transition line corresponds to the case with $r_{V}=1$, where the interaction part of the Hamiltonian has the full SU(2) spin rotation symmetry. Then for the 1D system, the SOC reduces this symmetry to the U(1) symmetry corresponding to the spin rotation around the special axis with unit vector $\hat{n}_{\alpha}=(\alpha_{D},0,\alpha_{R})/\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}$. Then one can always choose the spin quantization axis along the axis $\hat{n}_{\alpha}$ and the Hamiltonian thus becomes
@@
Figure 4: Properties of typical spin-triplet pair density wave (t-PDW) and mixed-parity (MP) states. a,b The band dispersion of t-PDW (a) and MP (b) states, where that of the t-PDW state is plotted in the folded Brillouin zone. c,d The energy spectrum obtained in a finite-size system with L=320, hosting four zero-energy MZMs denoted by blue and red dots. The parameters used are $r_{\rm so}$=1.4 and $r_{V}$=0.3 for the t-PDW state and $r_{\rm so}$=3.0 and $r_{V}$=0.3 for the MP state with $\mu=-0.4$ and all other parameters mentioned in the main text.
Real space modulations of the $d$-vector in t-PDW
We next study the obtained pairing order parameters of all the states in the phase diagram, which helps to visualize the spatial modulation of the t-PDW state. As mentioned above, the real space pairing order parameters can be compactly expressed by Eq. 1 as
@@
The real space $d$-vector is pure imaginary due to the time-reversal symmetry. The spatial modulation of a typical $d$-vector of the t-PDW state constructed from the pairing order parameters determined self-consistently at $r_{\rm so}=1.4$ and $r_{V}=0.3$ is shown schematically in Fig. 1, where the $d$-vector revolves around the origin in an elliptical orbit in 3D space within a period of the PDW state.
On the other hand, the KFF state at $r_{\rm so}=0$ and $r_{V}<1$ has only one finite order parameter $\Delta_{\parallel,Q}$, corresponding to the $d$-vector
@@
The band dispersion of both t-PDW and MP states are shown in Fig. 4a,b, both of which are fully gapped. In order to determine the topological properties of these states, we calculate the end states of finite-size systems in each phase. As shown in Fig. 4c,d, both phases host two pairs of Majorana zero modes (MZMs) at each end of the chain, indicating the nontrivial topological properties. To fully capture the topological properties of both states, we calculate their $Z_{2}$ topological invariant, $\cal N$}. While a standard method exists for calculating the $Z_{2}$ invariant for MP states [91], applying it to the translation symmetry breaking PDW states proves nontrivial. Therefore, we generalize this method to calculate the $Z_{2}$ invariant for the PDW state (see Methods). We apply this method to the t-PDW states and find that the $Z_{2}$ topological invariant ${\cal N}=-1$, consistent with the presence of topological boundary states, specifically the two pairs of MZMs shown in Fig. 4b. Since the time-reversal symmetry is preserved for both phases, they can be classified as time-reversal invariant (TRI) TSC. The nontrivial topological property of the t-PDW state is easy to understand since it is a pure $p$-wave spin-triplet pairing state with $d$-vector given by Eq. 8 and the spatial modulation of the $d$-vector does not change its topological property. In contrast, the MP state has both $s$-wave and $p$-wave pairing components and can be either a TSC or a topologically trivial superconductor [89]. For the chemical potential close to half-filling with $\mu=-0.4$, the MP state is always topological, which is labeled as MPTSC in the phase diagram of Fig. 3. However, a topological phase transition can occur when the chemical potential is tuned away from the half-filling, which is studied below.
@@
Finally, we study the effect of the chemical potential on the phase diagram. The calculation above is done for $\mu=-0.4$ which is around the half-filling of the chain. We find that if the chemical potential is tuned away from the half-filling, the region for the t-PDW phase shrinks as shown in Fig. 5a for $\mu=-1.4$. For $\mu$ close enough to the band top or bottom, the phase diagram changes dramatically, where the t-PDW state can only be realized for small values of $r_{V}$ and $r_{\rm so}$ which is shown in Fig. 5b with $\mu=-1.8$.
Figure 5: Phase diagrams away from half filling. The chemical potential is set as $\mu$=-1.4 (a) and $\mu=-1.8$ (b) with all the other parameters unchanged. The Kramers Fulde-Ferrell (KFF) states are realized along the $x$-axis, indicated by the bold blue lines. The narrow phase region shaded in grey and bounded by two nearby red lines in a corresponds to the time-reversal symmetry-breaking ss+ipp state, which emerges near the transition between two time-reversal invariant mixed-parity (MP) states without gap closing [92, 93]. The single red line in b represents the continuous phase transition between the topological superconductor (TSC) and trivial superconductor with mixed parity.
@@
where $t_{\alpha}=\sqrt{t^{2}+\alpha_{R}^{2}+\alpha_{D}^{2}}$, $\theta_{\alpha}=\arctan\left(\frac{\sqrt{\alpha_{R}^{2}+\alpha_{D}^{2}}}{t}\right)=\frac{Q}{2}$ and $n^{d}_{i\sigma}=d_{i\sigma}^{\dagger}d_{i\sigma}$. Since the SOC is gauged away in the transformed Hamiltonian, which respects the full spin rotation and inversion symmetry, the ground state of such a system is either the uniform $s$-wave pairing state or the $p$-wave state with its $d$-vector pointing in any direction, depending on the chemical potential. For $\mu$ around half filling, the Fermi points $k_{f}$ are round $\pm\frac{\pi}{2}$ where the pairing function for the $p$-wave state $\Delta_{p}(k)\propto\sin k$ gains more energy than that of the $s$-wave state $\Delta_{s}(k)\propto\cos k$, making the $p$-wave pairing state as the ground state. On the other hand, if $\mu$ is sufficiently tuned away from half-filling, the $s$-wave pairing will take over as $k_{f}$ approaches $\pm\pi$, where $\Delta_{s}(k)\propto\cos k$ gains more energy.